# SubZero: Random Subspace Zeroth-Order Optimization for Memory-Efficient LLM Fine-Tuning

### v1

@@ 4.1 Random Subspace Optimization for LLM Fine-Tuning
0:   parameter matrix in the $i$-th layer $\bm{W}_{i}\in\mathbb{R}^{m_{i}\times n_{i}},i=1,2,\dots,l$, loss $\mathcal{L}$, step budget $T$, perturbation scale $\varepsilon$, learning rate schedule $\{\eta^{t}\}$, subspace change frequency $T_{0}$, rank $r$.
@@
SubZero maintains just three small matrices per layer: a perturbation matrix $\bm{Z}_{i}\in\mathbb{R}^{r\times r}$, and two column-orthogonal matrices $\bm{U}_{i}\in\mathbb{R}^{m_{i}\times r}$ and $\bm{V}_{i}\in\mathbb{R}^{n_{i}\times r}$. This design enhances memory efficiency, as $r$ is generally much smaller than the size of the corresponding parameter matrix $\bm{W}_{i}\in\mathbb{R}^{m_{i}\times n_{i}}$ (i.e., $r\ll\min\{m_{i},n_{i}\}$). Morover, we employ in-place operations and per-layer parameter updates to estimate gradients and update parameters in parallel (see Appendix A.4). Consequently, SubZero uses significantly less GPU memory than previous methods while achieving similar or better performance. For example, fine-tuning OPT-1.3B (Zhang et al., 2022) on SST-2 (Socher et al., 2013b) using SGD (without momentum) in full-parameter scheme as shown in Table 3, SubZero requires only 6.8GB GPU memory, compared to 11.5GB for SGD, yielding a 1.6×1.6$\times$1.6 × improvement in memory efficiency, similar as illustrated in Fig. 1 (d).
Now we are ready to summarize the overall algorithm of SubZero in Algorithm 3. Each training iteration consists of three main steps. First, it obtains the projection matrices $\bm{U}_{i}^{t}$ and $\bm{V}_{i}^{t}$ using Algorithm 1 or directly adopts previous ones. Next, it computes the loss value difference $\rho$ with Eqn. (7) by applying Algorithm 2 to perturb all parameter matrices. Finally, SubZero updates all parameter matrices layer by layer, following Eqn. (9).
@@ 4.2 Integration into Fine-Tuning Schemes
We describe the integration of SubZero into full-parameter tuning (Aghajanyan et al., 2021) and three promient PEFT schemes: LoRA (Hu et al., 2022), prefix tuning (Li & Liang, 2021), and prompt tuning (Lester et al., 2021). Typically, SubZero can be easily incorporated into these fine-tuning schemes. However, it encounters a challenge with extremely non-square parameter matrices, which have far more rows than columns or vice versa. This issue is particularly prevalent in LoRA, which employs two low-rank matrices $\bm{A}_{i}\in\mathbb{R}^{m_{i}\times k}$ and $\bm{B}_{i}\in\mathbb{R}^{k\times n_{i}}$ to approximate a full matrix $\bm{W}_{i}^{\prime}\in\mathbb{R}^{m_{i}\times n_{i}}$, with $k\ll\min\{m_{i},n_{i}\}$, e.g., $k=8$ while $\min\{m_{i},n_{i}\}=2048$ used in (Zhang et al., 2024). Consequently, it is impossible to find a smaller rank $r\ll k$ to compute the gradient estimates of $\bm{A}_{i}$ and $\bm{B}_{i}$ using Eqn. (6), imposing a challenge when applying SubZero to this scenario.
@@ 5 Theoretical Analysis
In this section, we theoretically analyze why SubZero can reduce the variance of gradient estimates and accelerate convergence. Before the analysis, we first define some necessary notations:
${\bm{P}}={\rm bdiag}(\bm{V}_{1}\otimes\bm{U}_{1},\cdots,\bm{V}_{l}\otimes\bm{U}_{l})\!,\ {\bm{z}}\!=\![{\rm vec}({\bm{Z}}_{1})^{\mathsf{T}},\dots,{\rm vec}({\bm{Z}}_{l})^{\mathsf{T}}]^{\mathsf{T}}\!,\ \tilde{{\bm{z}}}\!=\![{\rm vec}(\tilde{{\bm{Z}}}_{1})^{\mathsf{T}},\dots,{\rm vec}(\tilde{{\bm{Z}}}_{l})^{\mathsf{T}}]^{\mathsf{T}}\!.$
@@
See its proof in Appendix A.5. Theorem 5 (a) provides the equivalent form (10) of our gradient estimation (8). By comparing this with the gradient estimation (5) in random subspace optimization (Nozawa et al., 2024; Roberts & Royer, 2023), we observe significant differences. First, our gradient estimation (10) accounts for the layer-wise structure of the network, requiring the projection matrix ${\bm{P}}$ to be block-diagonal, whereas in random subspace optimization, ${\bm{P}}$ is not. Additionally, our method introduces a layer-wise low-rank perturbation matrix, reflected by the block-diagonal structure of ${\bm{P}}$, with lazy updates to the column and row spaces defined by ${\bm{U}}_{i}$ and ${\bm{V}}_{i}$. In contrast, random subspace optimization simply requires ${\bm{P}}$ to be random. These distinctions highlight the key differences between our gradient estimation and existing methods in random subspace optimization.
Theorem 5 (b) guarantees that the distance $\Phi({\bm{x}})$ between the expected gradient estimate and the BP gradient in the subspace spanned by ${\bm{P}}$ is small. Moreover, by setting $\varepsilon=\frac{1}{q+4}$, the distance $\Phi({\bm{x}})$ is bounded by a constant $L_{2}/6$, independent of the parameter dimension $d$. This implies that the error in our gradient estimation does not scale with the extremely high parameter dimensions of LLMs, providing highly accurate gradient estimation—crucial for optimizing LLMs.
Next, we utilize a strictly convex quadratic loss to further analyze our gradient estimation in Eqn. (10). This choice is motivated by the fact that, after pretraining, the LLM parameters tend to converge toward a local minimum within a local basin, which can be well-approximated by a quadratic loss (Neyshabur et al., 2020).
@@
See its proof in Appendix A.5. Theorem 5 demonstrates several advantageous properties of our gradient estimation on the quadratic function. First, Eqn. (11) establishes the equivalence between the expected gradient estimation and the BP gradient within the subspace spanned by our projection matrix ${\bm{P}}$. Second, Eqn. (12) shows that, in this subspace, the variance of the gradient estimation scales linearly with the subspace dimension $q$. In contrast, the variance of gradient estimation (2) in MeZO depends linearly on the model’s parameter dimension $d$, which is significantly larger than $q$. Finally, Eqn. (13) reveals that the expected cosine similarity between our estimated gradient and the BP gradient within the subspace depends only on the subspace dimension $q\ll d$, indicating that our gradient estimation provides a highly accurate parameter update direction.
Building upon the above results, we can prove the convergence of our SubZero. {theorem}[] Let ${\bm{x}}^{*}$ = $\operatorname*{arg\,min}_{{\bm{x}}\in\mathbb{R}^{d}}f({\bm{x}})$, where $f\in C^{1,1}_{L_{1}}(\mathbb{R}^{d})$ and $f$ is convex. Suppose $\mathcal{E}_{k}=({\bm{e}}_{0},\cdots,{\bm{e}}_{k})$ where ${\bm{e}}_{k}\sim\mathcal{N}(0,{\bm{I}}_{q})$, $\eta=\frac{1}{4(q+4)L_{1}}$, $\phi_{0}=f({\bm{x}}_{0})$, $\phi_{k}=\mathbb{E}_{\mathcal{E}_{k-1}}\left[f\left(\bm{x}_{k}\right)\right],k\geq 1$ where $\{{\bm{x}}_{k}\}_{k>0}$ is the sequence generated by Algorithm 3. For a fixed $\bm{P}$, then after $N=\mathcal{O}(\frac{q}{\epsilon})$ training iterations, we have
$\frac{1}{N+1}\sum_{k=0}^{N}\left(\phi_{k}-f^{*}\right)\leq\epsilon.$
See its proof in Appendix A.5. Theorem 5 guarantees the convergence of our SubZero when the projection matrix $\bm{P}$ is fixed. Note, here we follow the approach of Galore (Zhao et al., 2024a), and assume a fixed projection matrix for simplicity. This convergence result can also be extended to cases where the projection matrix is lazily updated. Since lazy updates involve keeping the projection fixed over each periodic interval, we can prove convergence within each such period.
@@ 6 Experiments
In this section, we present comprehensive experiments to evaluate the effectiveness of SubZero. We conduct our experiments using medium-sized masked LLMs (RoBERTa-large (Liu et al., 2019)) and large-scale autoregressive LLMs (OPT-1.3B and 13B (Zhang et al., 2022), LLaMA2-7B (Touvron et al., 2023), and Mistral-7B (Jiang et al., 2023)). Our exploration covers full-parameter tuning (FT) (Aghajanyan et al., 2021) and three PEFT schemes: LoRA (Hu et al., 2022), prefix tuning (Li & Liang, 2021), and prompt tuning (Lester et al., 2021). For comparison, we include leading ZO methods, such as MeZO (Malladi et al., 2023) and S-MeZO (Liu et al., 2024), alongside inference-only memory-efficient baselines like zero-shot, in-context learning (ICL) (Brown et al., 2020), and linear probing (LP) (Kumar et al., 2022). We also use the FO optimizer SGD as a benchmark. Since appropriate prompts are critical for ZO optimization (Malladi et al., 2023; Zhang et al., 2024), all experiments incorporate prompt templates, which are detailed in Appendix A.1.
Table 2: Performance of fine-tuning OPT-13B on SuperGLUE with various experimental settings (with 1000 examples). AVG: average score of all tasks.
@@
SGD(FT) 94.9 82.3 85.7 78.4 65.3 65.8 74.2 90.0 82.4 88.0 35.5 76.6
Zero-shot 58.8 59.6 46.4 59.0 38.5 55.0 46.9 80.0 81.2 46.2 14.6 53.3
ICL 87.0 62.1 57.1 66.9 39.4 50.5 53.1 87.0 82.5 75.9 29.6 62.8
LP 93.4 68.6 67.9 59.3 63.5 60.2 63.5 55.0 27.1 3.7 11.1 52.1
MeZO(FT) 92.1 71.5 71.4 74.4 61.5 60.0 60.1 87.0 82.0 84.2 31.2 70.5
S-MeZO(FT) 92.3 76.9 75.0 76.5 61.1 58.2 63.3 87.0 71.2 77.9 31.9 70.1
SubZero(FT) 92.1 74.0 73.2 75.3 65.4 60.8 61.0 88.0 82.3 84.5 32.0 71.5
MeZO(LoRA) 92.2 74.4 69.6 75.2 64.4 59.7 58.2 87.0 82.0 82.9 31.0 70.6
S-MeZO(LoRA) 90.8 62.2 75.0 72.9 51.9 55.8 56.4 86.0 69.9 76.4 31.7 66.3
SubZero(LoRA) 93.8 75.5 71.4 76.1 65.4 60.3 60.3 89.0 81.9 83.7 31.3 71.7
@@
Cross entropy (SubZero) 92.9 54.0 84.7 77.1 84.5
@@
6.1 Performance with Different Experimental Settings
Following the settings in MeZO (Malladi et al., 2023), we evaluated SubZero using OPT-13B on the SuperGLUE benchmark (Wang et al., 2019), which covers a diverse range of tasks, including classification, multiple-choice, and generation, as outlined in Table 2. For each task, we randomly sampled 1000 examples for training, 500 for validation, and 1000 for testing. The ZO methods were applied to both full-parameter tuning (FT) and LoRA fine-tuning schemes, running for 20K steps.
Table 2 presents the key findings, highlighting the best-performing ZO method in bold. The results show that ZO techniques significantly outperform baseline approaches like zero-shot, in-context learning, and linear probing, underscoring their ability to enhance a pre-trained model’s performance on downstream tasks.
From Table 2, one can also observer that SubZero consistently surpasses MeZO across all tasks and fine-tuning methods. For instance, SubZero boosts MeZO’s accuracy from 61.1% to 65.4% on the WSC task (+4.3%) under FT, and from 58.2% to 60.3% on MultiRC using LoRA (+2.1%). S-MeZO demonstrated competitive performance on several classification tasks. However, SubZero outperformed S-MeZO in 6 out of 11 tasks with FT and 9 out of 11 tasks with LoRA. Additionally, SubZero’s average score across all tasks was higher than S-MeZO’s, which displayed inconsistent performance due to its selective parameter masking based on pre-determined thresholds—an approach that lacked robustness in practice. Moreover, S-MeZO’s performance in the LoRA scheme was particularly poor, highlighting the need for more adaptive sparse masking strategies.

### v3

@@ 4.1 Random Subspace Optimization for LLM Fine-Tuning
0:   parameter matrix in the $i$-th layer $\bm{W}_{i}\in\mathbb{R}^{m_{i}\times n_{i}},i=1,2,\dots,l$, loss $\mathcal{L}$, step budget $T$, perturbation scale $\varepsilon$, learning rate schedule $\{\eta^{t}\}$, subspace change frequency $F$, rank $r$.
@@
SubZero maintains just three small matrices per layer: a perturbation matrix $\bm{Z}_{i}\in\mathbb{R}^{r\times r}$, and two column-orthogonal matrices $\bm{U}_{i}\in\mathbb{R}^{m_{i}\times r}$ and $\bm{V}_{i}\in\mathbb{R}^{n_{i}\times r}$. This design enhances memory efficiency, as $r$ is generally much smaller than the size of the corresponding parameter matrix $\bm{W}_{i}\in\mathbb{R}^{m_{i}\times n_{i}}$ (i.e., $r\ll\min\{m_{i},n_{i}\}$). Moreover, we employ in-place operations and per-layer parameter updates to estimate gradients and update parameters in parallel (see Appendix 8.2). Consequently, SubZero uses significantly less GPU memory than previous methods while achieving similar or better performance. For example, fine-tuning OPT-1.3B [64] on SST-2 [51] using SGD (without momentum) in full-parameter scheme as shown in Table 5, SubZero requires only 6.8GB GPU memory, compared to 11.5GB for SGD, yielding a 1.6×1.6$\times$1.6 × improvement in memory efficiency, similar as illustrated in Fig. 1 (d).
Now we are ready to summarize the overall algorithm of SubZero in Algorithm 3. Each training step consists of three sequential phases. First, it obtains the projection matrices $\bm{U}_{i}^{t}$ and $\bm{V}_{i}^{t}$ using Algorithm 1 or directly adopts previous ones. Next, it computes the loss value difference $\rho$ with Eqn. (7) by applying Algorithm 2 to perturb all parameter matrices. Finally, SubZero updates all parameter matrices layer by layer, following Eqn. (9).
@@ 4.2 Integration into Fine-Tuning Schemes
We describe the integration of SubZero into full-parameter tuning [2] and three prominent PEFT schemes: LoRA [24], prefix tuning [36], and prompt tuning [33]. Typically, SubZero can be easily incorporated into these fine-tuning schemes. However, it encounters a challenge with extremely non-square parameter matrices, which have far more rows than columns or vice versa. This issue is particularly prevalent in LoRA, which employs two low-rank matrices $\bm{A}_{i}\in\mathbb{R}^{m_{i}\times k}$ and $\bm{B}_{i}\in\mathbb{R}^{k\times n_{i}}$ to approximate a full matrix $\bm{W}_{i}^{\prime}\in\mathbb{R}^{m_{i}\times n_{i}}$, with $k\ll\min\{m_{i},n_{i}\}$, e.g., $k=8$ while $\min\{m_{i},n_{i}\}=2048$ used in [65]. Consequently, it is impossible to find a smaller rank $r\ll k$ to compute the gradient estimates of $\bm{A}_{i}$ and $\bm{B}_{i}$ using Eqn. (6), imposing a challenge when applying SubZero to this scenario.
@@ 5 Theoretical Analysis
In this section, we theoretically analyze why SubZero can reduce the variance of gradient estimates and hence accelerate convergence. Before the analysis, we first define some necessary notations:
${\bm{P}}$ $={\rm bdiag}(\bm{V}_{1}\otimes\bm{U}_{1},\cdots,\bm{V}_{l}\otimes\bm{U}_{l})\!,$ (10)
${\bm{z}}$ $=[{\rm vec}({\bm{Z}}_{1})^{\mathsf{T}},\dots,{\rm vec}({\bm{Z}}_{l})^{\mathsf{T}}]^{\mathsf{T}}\!,$ (11)
$\tilde{{\bm{z}}}$ $=[{\rm vec}(\tilde{{\bm{Z}}}_{1})^{\mathsf{T}},\dots,{\rm vec}(\tilde{{\bm{Z}}}_{l})^{\mathsf{T}}]^{\mathsf{T}}\!.$ (12)
@@
See its proof in Appendix 8.6. Theorem 5 (a) provides the equivalent form (13) of our gradient estimation (8). By comparing this with the gradient estimation (5) in random subspace optimization [43, 46], we observe significant differences. First, our gradient estimation (13) accounts for the layer-wise structure of the network, requiring the projection matrix ${\bm{P}}$ to be block-diagonal, whereas in random subspace optimization, ${\bm{P}}$ is not. Additionally, our method introduces a layer-wise low-rank perturbation matrix, reflected by the block-diagonal structure of ${\bm{P}}$, with lazy updates to the column and row spaces defined by ${\bm{U}}_{i}$ and ${\bm{V}}_{i}$. In contrast, random subspace optimization simply requires ${\bm{P}}$ to be random. These distinctions highlight the key differences between our gradient estimation and existing methods in random subspace optimization.
Theorem 5 (b) guarantees that the distance $\Phi({\bm{x}})$ between the expected gradient estimate and the BP gradient in the subspace spanned by ${\bm{P}}$ is small. Moreover, by setting $\varepsilon=\frac{1}{q+4}$, the distance $\Phi({\bm{x}})$ is bounded by a constant $L_{2}/6$, independent of the parameter dimension $d$. This implies that the error in our gradient estimation does not scale with the extremely high parameter dimensions of LLMs, providing accurate gradient estimation—crucial for optimizing LLMs.
Next, we utilize a strictly convex quadratic loss to further analyze our gradient estimation in Eqn. (13). This choice is motivated by the fact that, after pretraining, the LLM parameters tend to converge toward a local minimum within a local basin, which can be well-approximated by a quadratic loss [42].
@@
See its proof in Appendix 8.6. Theorem 5 demonstrates several advantageous properties of our gradient estimation on the quadratic function. First, Eqn. (14) establishes the equivalence between the expected gradient estimation and the BP gradient within the subspace spanned by our projection matrix ${\bm{P}}$. Second, Eqn. (15) shows that, in this subspace, the variance of the gradient estimation scales linearly with the subspace dimension $q$. In contrast, the variance of gradient estimation (2) in MeZO depends linearly on the model’s parameter dimension $d$, which is significantly larger than $q$. Finally, Eqn. (16) reveals that the expected cosine similarity between our estimated gradient and the BP gradient within the subspace depends only on the subspace dimension $q\ll d$, indicating that our gradient estimation provides a highly accurate parameter update direction.
Building upon the above results, we can prove the convergence of our SubZero. {theorem}[] Let $f\in C^{1,1}_{L_{1}}(\mathbb{R}^{d})$ be a non-convex function bounded below by $f^{*}$. Suppose $\mathcal{E}_{k}=(\bm{z}_{0},\bm{z}_{1},\cdots,\bm{z}_{k})$ with $\bm{z}_{k}\sim\mathcal{N}(\bm{0},{\bm{I}}_{q})$, and let $\mathcal{P}_{j}=({\bm{P}}_{0},{\bm{P}}_{1},\cdots,{\bm{P}}_{j})$, where ${\bm{P}}_{j}$ is defined in (10) with a fixed update frequency $F$. Then, the sequence $\{{\bm{x}}_{k}\}_{k>0}$ generated by Algorithm 3 satisfies:
$\frac{1}{T}\sum_{k=0}^{T-1}\mathbb{E}_{\mathcal{E}_{k},\mathcal{P}_{\left\lfloor k/F\right\rfloor}}\left[\|\nabla f({\bm{x}}_{k})\|^{2}\right]\leq\epsilon$
with $T=\mathcal{O}\left(\frac{d}{\epsilon}\right)$ if the perturbation scale satisfies $\varepsilon\leq\mathcal{O}\left(\frac{\epsilon^{1/2}}{q^{3/2}d^{1/2}L_{1}^{3/2}}\right)$. Here $T=KF$, where $K$ denotes the total number of subspace updates. See its proof in Appendix 8.6. Theorem 5 guarantees the convergence of our SubZero when the projection matrix $\bm{P}$ is updated at a fixed frequency $F$.
@@ 6 Experiments
Table 3: Performance of fine-tuning OPT-13B on SuperGLUE with various experimental settings (with 1000 examples). AVG: average relative percentage difference with MeZO of all tasks.
@@
SGD(FT) 94.9 82.3 85.7 78.4 65.3 65.8 74.2 90.0 82.4 88.0 35.5 -
Zero-shot 58.8 59.6 46.4 59.0 38.5 55.0 46.9 80.0 81.2 46.2 14.6 -
ICL 87.0 62.1 57.1 66.9 39.4 50.5 53.1 87.0 82.5 75.9 29.6 -
LP 93.4 68.6 67.9 59.3 63.5 60.2 63.5 55.0 27.1 3.7 11.1 -
MeZO(FT) 92.1 71.5 71.4 74.4 61.5 60.0 60.1 87.0 82.0 84.2 31.2 0%
ZO-AdaMU(FT) 92.1 72.9 67.9 73.0 61.5 60.7 63.0 89.0 83.0 82.4 32.0 0.46%
S-MeZO(FT) 92.3 76.9 75.0 76.5 61.1 58.2 63.3 87.0 71.2 77.9 31.9 -0.10%
HiZOO(FT) 91.3 69.3 69.4 67.3 63.5 59.4 55.5 88.0 81.4 81.9 31.3 -2.15%
LOZO(FT) 92.9 73.6 71.4 70.7 63.5 60.2 60.3 87.0 81.7 84.5 30.7 0.10%
SubZero(FT) 92.1 74.0 73.2 75.3 65.4 60.8 61.0 88.0 82.3 84.5 32.0 1.89%
MeZO(LoRA) 92.2 74.4 69.6 75.2 64.4 59.7 58.2 87.0 82.0 82.9 31.0 0%
ZO-AdaMU(LoRA) 88.0 72.0 71.6 72.6 60.1 56.4 58.9 88.0 83.2 76.8 32.4 -1.78%
S-MeZO(LoRA) 90.8 62.2 75.0 72.9 51.9 55.8 56.4 86.0 69.9 76.4 31.7 -5.79%
HiZOO(LoRA) 90.6 67.5 69.6 70.5 63.5 60.2 60.2 87.0 81.9 83.8 31.2 -1.16%
SubZero(LoRA) 93.8 75.5 71.4 76.1 65.4 60.3 60.3 89.0 81.9 83.7 31.3 1.57%
(a) WIC: OPT-13B, FT
(b) SQuAD: OPT-13B, LoRA
(c) CB: LLaMA2-7B, Prompt
(d) MNLI: RoBERTa, FT, Non-diff
Figure 2: Training loss curves of MeZO and SubZero.
@@
Table 5: Orthogonal projection matrix.
Dataset Ortho. Accuracy
RTE 67.5
74.0
WSC 59.6
65.1
@@
Cross entropy (SubZero) 93.4 54.0 84.7 77.4 84.5
@@
In this section, we present comprehensive experiments to evaluate the effectiveness of SubZero. We conduct our experiments using medium-sized masked LLMs (RoBERTa-large [38]) and large-scale autoregressive LLMs (OPT-1.3B and 13B [64], LLaMA2-7B [55], and Mistral-7B [27]). Our exploration covers full-parameter tuning (FT) [2] and three PEFT schemes: LoRA [24], prefix tuning [36], and prompt tuning [33]. For comparison, we include leading ZO methods, such as MeZO [40], ZO-AdaMU [28], S-MeZO [39], HiZOO [68], and LOZO [9] alongside inference-only memory-efficient baselines like zero-shot, in-context learning (ICL) [7], and linear probing (LP) [32]. As the first and most popular ZO optimizer for LLM fine-tuning, MeZO is considered our primary competitor. We also use the FO optimizer SGD as a benchmark. Since appropriate prompts are critical for ZO optimization [40, 65], all experiments incorporate prompt templates. Since a larger batch size reduces the variance of the estimated gradients, all compared methods use a batch size of 16 unless otherwise specified. All experimental settings are detailed in Appendix 8.2-8.5.
6.1 Results under Different Experimental Settings
Following the settings in MeZO [40], we evaluated SubZero using OPT-13B on the SuperGLUE benchmark [57], which covers a diverse range of tasks, including classification, multiple-choice, and generation, as outlined in Table 3. The ZO methods were applied to both full-parameter tuning (FT) and LoRA fine-tuning schemes. The comparisons with vanilla LoRA and SGD with gradient accumulation are provided in Appendix 8.1.
Table 3 presents the key findings, highlighting the best-performing ZO method in bold. The results show that ZO techniques significantly outperform baseline approaches like zero-shot, in-context learning, and linear probing, underscoring their ability to enhance a pre-trained model’s performance on downstream tasks. From Table 3, one can also observer that MeZO, the first ZO optimizer for LLM fine-tuning, is highly competitive after carefully tuning its hyperparameters. Only ZO-AdaMU and LOZO, apart from SubZero, outperform MeZO in FT scheme. SubZero consistently surpasses MeZO across all tasks and fine-tuning schemes. In FT scheme, S-MeZO showed competitive performance on several classification tasks. However, its performance on ReCoRD remained unsatisfactory even after hyperparameter tuning. Excluding ReCoRD, SubZero still outperforms S-MeZO with 2.05% vs. 1.20% in FT scheme and 1.74% vs. -4.90% in LoRA scheme.